OpenStudy (rayman5766):

A patio in the shape of a semicircle has a diameter of 21 ft. What is the approximate area of the patio, rounded to the nearest square foot? Use 3.14 to approximate pi

OpenStudy (cggurumanjunath):

a = pi * d ^ 2 / ( 4 * 2 ) = ( 3 . 14 * 21 * 21 ) / 8

OpenStudy (rayman5766):

ok hold on let me get this

OpenStudy (cggurumanjunath):

3634.9425 ft ^ 2

OpenStudy (cggurumanjunath):

29079 . 54 / 8 ; 3.14 * 21 * 21 = 29079 . 54

OpenStudy (cggurumanjunath):

Area = 3635 ft ^ 2

OpenStudy (cggurumanjunath):

3635 is the final answer

OpenStudy (cggurumanjunath):

as they are asking to round to nearest square foot .

OpenStudy (rayman5766):

oh ok now I see thanks a lot

OpenStudy (cggurumanjunath):

:)